OpenStudy (anonymous):

Divide x to the 5 eighths power divided by x to the 1 fourth power.

OpenStudy (phi):

like this: \[ \frac{x^\frac{5}{8}}{x^\frac{1}{4}} \] if the problem was \[ \frac{x^2}{x^1} \\ \frac{x\cdot x}{x} = x = x^1\] or \[ \frac{x^4}{x^2} = \frac{x\cdot x \cdot x \cdot x}{ x \cdot x} = x\cdot x = x^2\] notice the "rule": the exponent of the numerator minus the exponent of the denominator. that rule still works even when the exponents are fractions!

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

got it